# Cosmic baryon census with fast radio bursts and gravitational waves

### v2

@@ II.3 Joint analysis of FRB and GW data
III Results and discussions
In this section, we first report the constraints on the baryon density as our main results in Sec. III.1, and then extend the analysis to assess the impacts of FRB and GW population modeling assumptions in Secs. III.2 and III.3, respectively.
III.1 Main results
First, we report the main results for the baryon-density estimate in the first block of Table 1. The posteriors for $\Omega_{\rm b}$~and the corresponding credible intervals are shown in Fig. 1. Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level. Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs. Remarkably, this late-Universe estimate exhibits concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level, linking the gap between the early and late Universe. Note that a similarly precise constraint is simultaneously obtained for the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{kms^{-1}~Mpc^{-1}}$, primarily driven by the GW dataset alone, which gives $H_{0}=72.6^{+5.2}_{-8.9}~\mathrm{kms^{-1}~Mpc^{-1}}$. More data is required to arbitrate the $H_{0}$ tension. A detailed discussion on $H_{0}$ and DM parameters is provided in Appendix A.
We compare our $\Omega_{\rm b}$~results with those from previous works. Macquart20 and Yang22 constrained $\Omega_{\rm b}$~using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Fig. 1, respectively. The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior. Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Fig. 1. This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates. Although the induced $\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses. Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder. This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang:2023gye.
We also assess the dependence on the prior for $\Omega_{\rm m}$. Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1, as shown in the third row of Table 1. Compared with the fiducial case ($\Omega_{\rm b}=0.0488\pm 0.0064$), the slightly lower value of $\Omega_{\rm b}$ is driven by the biased constraint on $\Omega_{\rm m}$ under the broad prior, $\Omega_{\rm m}=0.225^{+0.060}_{-0.180}$, which propagates through the positive correlation between $\Omega_{\rm b}$~and $\Omega_{\rm m}$. Therefore, we caveat that our present analysis moderately depends on the $\Omega_{\rm m}$ prior, which can be improved with larger future FRB and GW samples to better constrain $\Omega_{\rm m}$.
III.2 Impact of FRB model parameters
Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (indigo) data. The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies. Note that $H_{0}$ is in units of km s-1  Mpc-1.
Next, we assess the impact of FRB population modeling on the baryon census, as listed in the second block of Table 1. In our fiducial analysis in Sec. III.1, no specific FRB population model is imposed and all DM parameters are left free to avoid potential bias. To explore how such modeling affects the $\Omega_{\rm b}$ inference, we apply several commonly used FRB population models and evaluate their potential to improve the precision of $\Omega_{\rm b}$. FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$. To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts. Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) Zhang:2020mgq; Zhang:2020xoc; and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB. The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Fig. 1. Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias. However, such agreement at the current level does not ensure the absence of systematic errors when applied to larger future FRB and GW samples, in which systematic errors may become dominant. Therefore, a detailed assessment of how the DM parameters affect the inferred baryon density is of great importance.
To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset. The lower row of Fig. 2 shows the marginalized posteriors of $\Omega_{\rm b}$~and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs (r=0.73r=-0.73), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined. While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements. Therefore, the $\Omega_{\rm b}$ constraint from FRB + GW is robust against the $H_{0}$-related bias and free from assumptions affected by the $H_{0}$ tension.
On the other hand, the correlations between $\Omega_{\rm b}$~and DM parameters are moderate. These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies James:2022dcx, and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$. In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are 0.28-0.28, $0.40$, 0.40-0.40, and $0.17$, respectively. For comparison, the upper row of Fig. 2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~\mathrm{km~s^{-1}~Mpc^{-1}})$. This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone Macquart:2020lln; connor2025gas. In this case, the corresponding correlations are notably stronger: 0.41-0.41, $0.57$, 0.62-0.62, and $0.27$, respectively. This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
We conclude that the impact of FRB population modeling on the inferred $\Omega_{\rm b}$ is negligible within our analysis. The results also show that, within our FRB + GW framework, incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters. This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
III.3 Impact of GW model parameters
@@
Figure 3: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the typical GW population parameter $\mu_{g}$ using FRB + GW data. Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalised over (gray). The dotted lines indicate the median values of each posterior, except for $\mu_{g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{g}=32.27$ adopted following Ref. LIGOScientific:2021aug. Note that $H_{0}$ is in units of km s-1  Mpc-1.
@@
Last, we assess the impact of GW population modeling on our baryon census, as listed in the third block of Table 1. In our fiducial analysis in Sec. III.1, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference LIGOScientific:2021aug. To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them. We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$. The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains consistent with the CMB + BBN determination at the $1\sigma$ level.
@@
To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model. This parameter is moderately correlated with the inferred $H_{0}$ from GW observations LIGOScientific:2021aug, making it suitable for accounting for the impact of population modeling. As shown in Fig. 3, $\mu_{g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$. Consequently, compared with the prior value $\mu_{g}=32.27$ adopted in the GW–parameter-fixed (indigo) case, the lower inferred value of $\mu_{g}=31.5$ in the GW–parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$. Thus, within our FRB + GW framework, the inferred $\Omega_{\rm b}$ is more sensitive to GW-related parameters than to FRB-related ones (see Sec. III.2). It is because $H_{0}$, the essential parameter to infer $\Omega_{\rm b}$ in our framework, is more strongly degenerate with the GW population parameters than with the FRB-related parameters.
We conclude that, to obtain an unbiased and robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$. Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
IV Conclusion
@@
In this paper, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach. Most previous baryon censuses assumed a fixed value of $H_{0}$ or adopted an external prior, which could bias the inference in the presence of the Hubble tension. If FRB DMs are instead used to trace the baryons, additional parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$. GWs independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy. This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem. We also obtain a simultaneous constraint on the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{km\,s^{-1}\,Mpc^{-1}}$, primarily driven by the GW data. Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively. These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced $\Omega_{\rm b}$~inconsistency” remains within $1\sigma$ uncertainty in our sample. (ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively. These two results agree with that of FRB + GW at the $1\sigma$ level, indicating that simulation-based priors can improve current constraints without introducing significant bias. However, in both (i) and (ii), systematic effects from external assumptions (related to $H_{0}$, $f_{\rm d}$, and empirical DM model parameters) may become non-negligible when applied to larger future FRB and GW samples.
Hence, we assess the associated systematic uncertainties through a detailed correlation analysis. On the one hand, the dominant uncertainty is the degeneracy between $\Omega_{\rm b}$ and $H_{0}$, implying the importance of carefully addressing the $H_{0}$-related bias. Unlike prior-based analyses that rely on assumed values of $H_{0}$, the joint FRB + GW analysis avoids such bias by uniquely inferring $H_{0}$ from the absolute distances provided by GWs, independent of both CMB and local distance ladder. On the other hand, the correlations between $\Omega_{\rm b}$ and DM parameters—treated as nuisance parameters—are moderate, and notably weaker than those of $\Omega_{\rm b}h_{70}$, making the direct inference of $\Omega_{\rm b}$ more robust against the uncertainties induced by the nuisance parameters. Therefore, combining FRB and GW data not only avoids the bias induced by the $H_{0}$ tension, but also mitigates model-dependent uncertainties.
@@
We caveat that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model. Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN values at the $1\sigma$ level. Using larger samples of localized FRBs and GW standard sirens, future analyses would jointly constrain $\Omega_{\rm m}$, GW population parameters, and $\Omega_{\rm b}$ more robustly, thereby reducing this dependence.
@@
Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts. With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers masui2023pinpointing and DSA-110 Law:2023ibd, their synergy will soon emerge as a powerful low-redshift cosmological probe.
V Acknowledgments

### v3

@@ Introduction.
Results.
Figure 1: Posteriors for $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo). The joint constraint is $\Omega_{\rm b}=0.0488\pm 0.0064$, obtained by marginalized over all DM parameters. Errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 Aghanim et al. (2020) and SH0ES Riess et al. (2022), respectively; Shaded bands in different colors show the results from CMB + BBN Cooke et al. (2018), Macquart20, and Yang22 Yang et al. (2022). Note that Yang22 adopts the prior on DM model parameters from the TNG simulation Zhang et al. (2020, 2021), a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$. All results are at the $1\sigma$ level.
@@ Results.
Figure 2: Posteriors ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the FRB DM parameter $f_{\rm d}$, $F$, $\sigma_{\rm host}$, $\mu_{\rm host}$, and the typical GW population parameter $\mu_{\rm g}$ from FRB + GW data. Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalized over (gray). The dotted lines indicate the median values of each posterior, except for $\mu_{\rm g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{\rm g}=32.27~M_{\odot}$ adopted following Ref. Abbott et al. (2023b).
We report the main results for the baryon-density estimate in the first block of Table 1. The posteriors for $\Omega_{\rm b}$~ and the corresponding credible intervals are shown in Fig. 1. Figure 2 shows the joint posterior distributions of the cosmological and DM parameters. Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level. Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs. The $\Omega_{\mathrm{b}}$-$H_{0}$ panel in Fig. 2 shows that this correlation is markedly stronger than those between $\Omega_{\mathrm{b}}$ and the other parameters. This late-Universe estimate exhibits remarkable concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level. For the FRB DM parameters, we obtain $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$, corresponding to a median rest-frame host DM of $108\pm 20~\mathrm{pc~cm^{-3}}$ and a mean host DM of $168^{+57}_{-40}~\mathrm{pc~cm^{-3}}$, consistent with previous estimates Yang et al. (2025); Connor et al. (2025).
We compare our $\Omega_{\rm b}$~results with those from previous works. Macquart20 and Yang22 constrained $\Omega_{\rm b}$~using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Fig. 1, respectively. The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior. Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Fig. 1. Obviously, the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates. Although the induced $\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses. Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and the local distance ladder. This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang et al. (2023).
@@
We assess the impact of GW population modeling on our baryon census. In our main analysis, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference Abbott et al. (2023b). To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them. We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$. The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains broadly consistent with the CMB + BBN determination at the $1\sigma$ level.
@@
To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{\rm g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model. This parameter is closely correlated with the inferred $H_{0}$ from GW observations Abbott et al. (2023b), making it suitable for accounting for the impact of population modeling. As shown in Fig. 2, $\mu_{\rm g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$. Consequently, compared with the prior value $\mu_{\rm g}=32.27~M_{\odot}$ adopted in the GW-parameter-fixed (indigo) case, the lower inferred value of $\mu_{\rm g}=31.5~M_{\odot}$ in the GW-parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$. Thus, to obtain a robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$. Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
We also assess the dependence on the prior for $\Omega_{\rm m}$. Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1. Compared with the fiducial case ($\Omega_{\rm b}=0.0488\pm 0.0064$), the slightly lower value of $\Omega_{\rm b}$ is driven by the biased constraint on $\Omega_{\rm m}$ under the broad prior, $\Omega_{\rm m}=0.225^{+0.060}_{-0.180}$, which propagates through the positive correlation between $\Omega_{\rm b}$~and $\Omega_{\rm m}$. Therefore, we caveat that our present analysis moderately depends on the $\Omega_{\rm m}$ prior, which can be improved with larger future FRB and GW samples to better constrain $\Omega_{\rm m}$.
Figure 3: Same as Fig. 1 but for posteriors for $H_{0}$. The joint FRB + GW constraint is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior Cooke et al. (2018)) and GW data alone, respectively. The green, yellow, red, and gray shaded bands denote the results from Planck18 Aghanim et al. (2020), SH0ES Riess et al. (2022), Refs. Wu et al. (2022) (Wu22), and James et al. (2022) (James22), respectively. Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN Cooke et al. (2018) and CMB Aghanim et al. (2020), respectively. All results are at $1\sigma$ uncertainty.
Last, the corresponding constraints on $H_{0}$ are also listed in Table 1, with the posteriors and credible intervals shown in Fig. 3. A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies. Concretely, our FRB-only measurement gives $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, while including GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. This improvement is primarily driven by the GW dataset itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1. Previous FRB studies commonly inferred $H_{0}$ by fixing the baryon density to early-Universe values. For example, Refs. Wu et al. (2022) and James et al. (2022) inferred $H_{0}$ from different FRB samples by adopting external priors on $\Omega_{\rm b}h^{2}$, namely $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, as illustrated by the red and gray bands in Fig. 3. Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1 from FRB + $\Omega_{\rm b}h^{2}$, with a precision comparable to that of FRB + GW. Unlike these prior-based methods, the FRB + GW constraint provides a late-Universe determination of $H_{0}$. More data is required to offer further insight into arbitrating the $H_{0}$ tension.
Conclusion.
@@
In this Letter, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach. Previous baryon censuses often fixed $H_{0}$ or used external priors, which could bias the inference under the Hubble tension. If FRB DMs are employed to trace the baryons, serious parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$. GWs independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy. This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem. We also obtain a simultaneous constraint on the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{km\,s^{-1}\,Mpc^{-1}}$, primarily driven by the GW data. Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
@@
We note that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model. Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN determinations at the $1\sigma$ level. Using larger samples of localized FRBs and GW standard sirens, future analyses would jointly constrain $\Omega_{\rm m}$, GW population parameters, and $\Omega_{\rm b}$ more robustly, thereby reducing this dependence.
@@
Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts. With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers Masui (2023) and DSA-110 Law et al. (2024), their synergy will soon emerge as a powerful low-redshift cosmological probe.
Acknowledgments.